OpenStudy (anonymous):

writing equations here. your salary after a 5% salary increase if your salary before the increase was s. is the answer s-.05+s? how do i include the "if" in the equation? Please explain this to me

OpenStudy (anonymous):

I dont think that you have to do anything with that if s*1.05=x where x is your new salary or s+s*0.05=x (but this is the same)

OpenStudy (anonymous):

why 1.05 when the problem stated 5%?